# Where Quantum Complexity Helps Classical Complexity: A Survey

### v2

@@ 5.16. Others
Problem 89 ().
@@
Problem 90 ().
The problem of maximum cardinality matching can be optimized using 2000Q, a QA hardware from D-Wave. A better version of the QUBO formulation was used for this purpose. As part of their energy gap scale investigations, graphs with up to $22$ edges were fabricated. It is worth exploring whether this analysis can be extended to a $G_{2}$ graph. Doing so would provide direct insights into the original Graph family and enable verification of its scaling behavior with the maximum match problem for the $G_{n}$ graph family.
Problem 91 ().
@@
Problem 94 ().
The scholarly work cited in (anshu2023nlts, ) posits that quantum local testability is not a necessary condition for the construction of the work. The property, as established in the aforementioned work (Property 1), suffices for clustering the classical distributions of low-energy states. However, it is important to note that the property in question is weaker than local testability. Extant literature establishes that local testability implies NLTS. It is worthwhile to investigate the implications codes with Property 1 have for the quantum PCP conjecture (aharonov2013guest, ).
Problem 95 ().
The study discussed in (anshu2023nlts, ) has successfully established a property referred to as Property 1. Upon analyzing from the perspective of chain complexes, Property 1 seems to exhibit a relationship with the small-set boundary and co-boundary expansion as specified under Definition 1.2 of (hopkins2022explicit, ). This relationship may introduce a classical analog of the NLTS property. Furthermore, (hopkins2022explicit, ) presents a construction technique for classical Hamiltonians that are hard to approximate. The construction technique leverages the sum-of-squares hierarchy.
Problem 96 ().
Consider analyzing the subgraph isomorphism problems on a quantum gate-based computer (mariella2023quantum, ). Is it possible to generalize the problem (e.g., digraph isomorphism, commonly induced subgraphs, labeled edges and nodes), improve the compilation of the resulting circuit, improve classical solvers, and a comparative study of different Ansatze?
Problem 97 ().
Consider basketing, a post-selection technique (harris2023scalable, ). This allows to address time-dependent noise behaviors and reduces overhead. Is it possible to explore another case, in which the client suffers from state-preparation noise? In this case, one may consider more general classes of computation.
Problem 98 ().
@@
Problem 99 ().
[…]
Problem 100 ().
@@
Problem 105 ().
Consider a post-selection technique named basketing. This allows to address time-dependent noise behaviors and reduces overhead (harris2023scalable, ). Is there another case where the client suffers from state-preparation noise? In this case, one may consider more general classes of computation.
@@ 6. Conclusion
    Quantum Supremacy: The question of whether quantum computers can surpass classical computers (quantum supremacy) is a common one, and is explored in problems 1950, and 104. The concept of quantum supremacy is multifaceted and raises several questions: What precisely is quantum supremacy? Is it restricted to certain types of problems or is it all-encompassing? How can we demonstrate that quantum computers can solve problems that are practically infeasible tasks for classical computers?
@@
    Quantum Error Correction: The management and control of qubits presents a formidable challenge in the context of quantum computing, given the system’s susceptibility to noise and errors during computation. The detection and correction of these errors is an essential task to enhance the reliability and scalability of such computing systems. Collectively, problems 612263797, and 105 represent exemplars of the challenges that arise in this arena, and underscore the importance of developing effective strategies for addressing them.
@@
    Quantum Algorithms: To employ quantum hardware instead of classical alternatives, it is necessary to develop algorithms that are congruent with these devices. To fully harness the supremacy of quantum computing, it is imperative that we design algorithms that are more efficient than their classical counterparts. Examples of such algorithms are presented in problems:  615161723384967, and 103.
@@
    Hybrid Quantum-Classical Computing: The integration of classical and quantum computers can be a potent approach to solving intricate problems. This is typically achieved by allocating specific tasks to each type of computer. Two exemplary research papers that utilized hybrid computation in their work are Problems 14 and 39. In light of the complex nature of contemporary problems, the amalgamation of classical and quantum computing has emerged as a promising avenue for addressing hitherto intractable computational challenges. The judicious allocation of tasks between classical and quantum computers can lead to significant improvements in computational efficiency and accuracy.
@@
    Quantum Simulation: Quantum systems are characterized by their complexity and multi-dimensional nature, rendering their simulation on classical computers challenging. Nonetheless, the simulation of a quantum system utilizing another quantum system may offer a more efficient alternative. This avenue of approach could potentially unlock novel possibilities in the fields of materials science, as well as chemistry. Two illustrative examples of problems germane to this subject matter are problems 32 and 33.
@@
    Quantum Cryptography: Quantum computation has the potential to significantly impact cryptography due to the unique properties of quantum mechanics, such as entanglement. However, the extent to which quantum computing can enhance security protocols, such as quantum key distribution and secure communication, remains unclear. Notwithstanding, numerous problems in quantum cryptography have been identified, as enumerated by problems 4445545585, and 86.
@@
    Scalability: Quantum technologies have the potential to be revolutionary, but their scalability remains a challenge. Currently, controlling quantum systems and reducing noise is not fully feasible, resulting in a limited number of qubits in quantum computers. As problems 25814, and 35 highlight, building larger and more stable quantum systems is essential for their advancement.
@@
    Quantum Software and Programming Languages: As previously stated, quantum computers operate using distinctive algorithms. Consequently, the integration of quantum computers into practical applications necessitates the development of user-friendly programming languages and tools. It is highly desirable to have a quantum processor capable of executing any program to obtain the solution. However, the development of such a precisely programmable processor appears to be an insurmountable challenge. This issue is discussed in depth in Problem 9.
@@
    Quantum Complexity Theory: Quantum computational complexity classes have a unique definition that differs from the classical one. The relationship between these classes and classical complexity classes is unknown. Problems 43 to 77 are related to this topic. Problems: 43596365666869707375, and 77 are about this subject.
@@ 7.2. Competing Interests
7.3. Authors’ contributions
The idea was for the first author. All of the other authors are sorted based on their contributions.

### v4

@@ 6.16. Others
Problem 88.
Is it possible to find a fixed-parameter algorithm for Connected Planer
-Deletion
Connected Planer
-Deletion
@@
with a single-exponential dependence on $k$ s.t. $k\in\mathbb{N}$? The existing work suggests a factor-$(2+\varepsilon)$ parameterized approximation algorithm for the problem, with the running time of $2^{\mathcal{O}(k\log k)}n^{\mathcal{O}(1)}$ s.t. $10<\varepsilon<$ 1.
Problem 89.
The problem of maximum cardinality matching can be optimized using 2000Q, a QA hardware from D-Wave. A better version of the QUBO formulation was used for this purpose. As part of their energy gap scale investigations, graphs with up to $22$ edges were fabricated. It is worth exploring whether this analysis can be extended to a $G_{2}$ graph. Doing so would provide direct insights into the original Graph family and enable verification of its scaling behaviour with the maximum match problem for the $G_{n}$ graph family.
Problem 90.
@@
Problem 93.
The work in (anshu2023nlts, ) posits that quantum local testability is not a necessary condition for the construction of the work. The property, as established in the aforementioned work (Property 1), suffices for clustering the classical distributions of low-energy states. However, it is important to note that the property in question is weaker than local testability. Extant literature establishes that local testability implies NLTS. It is worthwhile to investigate the implications codes with Property 1 have for the quantum PCP conjecture (aharonov2013guest, ).
Problem 94.
The study in (anshu2023nlts, ) has successfully established a property referred to as Property 1. Upon analyzing from the perspective of chain complexes, Property 1 seems to exhibit a relationship with the small-set boundary and co-boundary expansion as specified under Definition 1.2 of (hopkins2022explicit, ). This relationship may introduce a classical analog of the NLTS property. Furthermore, (hopkins2022explicit, ) presents a construction technique for classical Hamiltonians that are hard to approximate. The construction technique leverages the sum-of-squares hierarchy.
Problem 95.
Consider analyzing the subgraph isomorphism problems on a quantum gate-based computer (mariella2023quantum, ). Is it possible to generalize the problem (e.g., digraph isomorphism, commonly induced subgraphs, labelled edges and nodes), improve the compilation of the resulting circuit, improve classical solvers, and a comparative study of different Ansatze?
Problem 96.
Consider basketing, a post-selection technique (harris2023scalable, ). This allows for the addressing of time-dependent noise behaviours and reduces overhead. Is exploring another case where the client suffers from state-preparation noise possible? In this case, one may consider more general classes of computation.
Problem 97.
@@
Problem 98.
There already exists a quantum oracle $\mathcal{O}$ s.t. $\text{BQP}^{\mathcal{O}}=\text{QMA}^{\mathcal{O}}$ and pseudorandom unitary transformations (PRUs) (and hence pseudorandom states (PRSs)) exist relative to $\mathcal{O}$. Do either PRUs or PRSs exist relative to $\mathcal{O}$ for a classic oracle?
Problem 99.
@@ 7. Conclusion
    Quantum Supremacy: The question of whether quantum computers can surpass classical computers (quantum supremacy) is a common one, and is explored in problems 1949, and 103. The concept of quantum supremacy is multifaceted and raises several questions: What precisely is quantum supremacy? Is it restricted to certain types of problems or is it all-encompassing? How can we demonstrate that quantum computers can solve problems that are practically infeasible tasks for classical computers?
@@
    Quantum Error Correction: The management and control of qubits presents a formidable challenge in the context of quantum computing, given the system’s susceptibility to noise and errors during computation. The detection and correction of these errors is an essential task to enhance the reliability and scalability of such computing systems. Collectively, problems 6122636, and  96 represent exemplars of the challenges that arise in this arena, and underscore the importance of developing effective strategies for addressing them.
@@
    Quantum Algorithms: To employ quantum hardware instead of classical alternatives, it is necessary to develop algorithms that are congruent with these devices. To fully harness the supremacy of quantum computing, it is imperative that we design algorithms that are more efficient than their classical counterparts. Examples of such algorithms are presented in problems:  615161723374866, and 102.
@@
    Hybrid Quantum-Classical Computing: The integration of classical and quantum computers can be a potent approach to solving intricate problems. This is typically achieved by allocating specific tasks to each type of computer. Two exemplary research papers that utilized hybrid computation in their work are Problems 14 and 38. In light of the complex nature of contemporary problems, the amalgamation of classical and quantum computing has emerged as a promising avenue for addressing hitherto intractable computational challenges. The judicious allocation of tasks between classical and quantum computers can lead to significant improvements in computational efficiency and accuracy.
@@
    Quantum Simulation: Quantum systems are characterized by their complexity and multi-dimensional nature, rendering their simulation on classical computers challenging. Nonetheless, the simulation of a quantum system utilizing another quantum system may offer a more efficient alternative. This avenue of approach could potentially unlock novel possibilities in the fields of materials science, as well as chemistry. Two illustrative examples of problems germane to this subject matter are problems 31 and 32.
@@
    Quantum Cryptography: Quantum computation has the potential to significantly impact cryptography due to the unique properties of quantum mechanics, such as entanglement. However, the extent to which quantum computing can enhance security protocols, such as quantum key distribution and secure communication, remains unclear. Notwithstanding, numerous problems in quantum cryptography have been identified, as enumerated by problems 4344535484, and 85.
@@
    Scalability: Quantum technologies have the potential to be revolutionary, but their scalability remains a challenge. Currently, controlling quantum systems and reducing noise is not fully feasible, resulting in a limited number of qubits in quantum computers. As problems 25814, and 34 highlight, building larger and more stable quantum systems is essential for their advancement.
@@
    Quantum Software & Programming Languages: As previously stated, quantum computers operate using distinctive algorithms. Consequently, the integration of quantum computers into practical applications necessitates the development of user-friendly programming languages and tools. It is highly desirable to have a quantum processor capable of executing any program to obtain the solution. However, the development of such a precisely programmable processor appears to be an insurmountable challenge. This issue is discussed in depth in Problem 9.
@@
    Quantum Complexity Theory: Quantum computational complexity classes have a unique definition that differs from the classical one. The relationship between these classes and classical complexity classes is unknown. Problems 42 to 76 are related to this topic. Problems: 42586264656768697274, and 76 are about this subject.